3.9GHz status
Friday, 3 October 2008 -
09:00
Monday, 29 September 2008
Tuesday, 30 September 2008
Wednesday, 1 October 2008
Thursday, 2 October 2008
Friday, 3 October 2008
09:00
Status of cavities 4, 5,6, 7 & 8
Status of cavities 4, 5,6, 7 & 8
09:00 - 09:20
Room: The Hermitage
09:20
Cavity 3 horizontal test results
Cavity 3 horizontal test results
09:20 - 09:40
Room: The Hermitage
09:40
MP9 cavity string preparation (status of travelers, required tooling, overall plans, etc.)
MP9 cavity string preparation (status of travelers, required tooling, overall plans, etc.)
09:40 - 10:00
Room: The Hermitage
10:00
ORC Documentation
-
H. Carter
ORC Documentation
H. Carter
10:00 - 10:20
Room: The Hermitage
10:20
Piping TIG welding
Piping TIG welding
10:20 - 10:40
Room: The Hermitage
10:40
Other business
Other business
10:40 - 11:00
Room: The Hermitage